64X=XSQUARE

I TIRED DIVIDED IT BUT IT DIDN'T WORK AND THEN I TIRED TO FIND THE SQUARE ROOT TO

Found 2 solutions by AnlytcPhil, stanbon:
Answer by AnlytcPhil(1806)

64X=XSQUARE 

I TIRED DIVIDED IT BUT IT DIDN'T WORK AND 
THEN I TIRED TO FIND THE SQUARE ROOT TO

-------------------------------------------

          64x = x²

Subtract x² from both sides:

     64x - x² = 0

Factor out x on the left

    x(64 - x) = 0

Set each of the factors, x and (64 - x) equal to 0.

x = 0 is one solution

Setting 64 - x = 0
            -x = -64
             x = 64

so 64 is the other solution.

Edwin

Answer by stanbon(75887)

x^2=64x
x^2-64x=0
x(x-64)=0
x=0 or x=64
Cheers,
Stan H.